# Align & Invert: Solving Inverse Problems with Diffusion and Flow-based Models via Representational Alignment

### v1

@@ 6 Conclusions
In this work, we introduced representation alignment as a principled framework for solving inverse problems with diffusion- and flow-based generative models. By aligning internal model representations with features from a pretrained DINOv2 encoder, our method provides a strong semantic prior that guides the reconstruction process toward perceptually meaningful solutions. We further offered theoretical insights into this mechanism, showing that REPA regularization corresponds to minimizing a divergence in the DINOv2 feature space and reduces the discrepancy between the model’s internal representations and those of the clean image. These results elucidate why feature-space alignment can enhance perceptual fidelity in inverse problem settings. Overall, our study demonstrates that representation alignment constitutes a flexible and effective tool for improving generative inverse problem solvers, yielding benefits at inference time and integrate seamlessly with a range of diffusion- and flow-based methods.
Appendix A Appendix
@@ A.1 Proofs
Proposition 0.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$ (Section 3), with $\bar{x}\sim p_{\bar{X}}$. Define the REPA alignment as
@@
and the empirical mean DINOv2 feature embedding
$\mu_{f}(x):=\frac{1}{N}\sum_{n=1}^{N}f_{\mathrm{DINOv2}}^{[n]}(x).$
Then the expected REPA alignment satisfies
@@
$),\frac{1}{N}\sum_{n=1}^{N}\big\|f_{\text{DINOv2}}^{[n]}(\bar{x})-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big\|_{2}^{2}=2\big(1-\text{REPA}(\bar{x},\hat{x}$$\big)$,
@@
Proposition 1.
Consider an inverse problem in which an observation $y\in\mathbb{R}^{m}$ is generated from a ground-truth image $x\sim p_{X}$, and let $\bar{x}\sim p_{\bar{X}}$ be the proxy approximation used during REPA (Section 3). Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at timestep $t$ produced by Algorithm 1 given $y$, and let $z_{t}^{\mathrm{(REPA)}}$ denote the diffusion state after applying the Repa regularization update. Let $z^{*}$ denote the latent variable associated with the clean image $x$, and define
$h_{t}=G_{2}(z_{t},t),\qquad h_{t}^{\mathrm{(REPA)}}=G_{2}(z_{t}^{\mathrm{(REPA)}},t),\qquad h^{*}=G_{2}(z^{*},0),$
where $G_{2}$ denotes the feature extractor returning the patch-level representation in layer $\ell$ of the model. Assume that the projection head $g_{\phi}$ is linear and that all DINOv2 patch embeddings are $\ell_{2}$–normalized. Then there exists a regularization weight $\lambda>0$ (Algorithm 1) such that the REPA update induces an approximate contraction, i.e., there exists a constant $C_{1}\in(0,1)$ and a constant $C_{2}\geq 0$ satisfying
$\|h_{t}^{\mathrm{(REPA)}}-h^{*}\|_{F}\;\leq\;C_{1}\,\|h_{t}-h^{*}\|_{F}+C_{2}\!\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right).$
Proof. We first rewrite all quantities in vectorized form. Let $\mathrm{vec}(\cdot)$ denote the operator that stacks the rows of a matrix into a single vector. Thus, for any matrix $H\in\mathbb{R}^{k_{1}\times k_{2}}$ we define
@@
Applying this to our setting, for the DINOv2 features $f_{\mathrm{DINOv2}}(x)\in\mathbb{R}^{N\times D_{1}}$ we write
$\widetilde{f}(x):=\mathrm{vec}\!\left(f_{\mathrm{DINOv2}}(x)\right)\in\mathbb{R}^{ND_{1}}.$
@@
For the diffusion-model feature extractor $G_{2}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{N\times D_{2}}$ we define its vectorized form
$\widetilde{G}_{2}(z_{t}):=\mathrm{vec}\!\left(G_{2}(z_{t},t)\right)\in\mathbb{R}^{ND_{2}}.$
Accordingly, we set
$\widetilde{h}_{t}:=\widetilde{G}_{2}(z_{t}),\qquad\widetilde{h}_{t}^{(\mathrm{repa})}:=\widetilde{G}_{2}(z_{t}^{(\mathrm{repa})}),\qquad\widetilde{h}^{*}:=\widetilde{G}_{2}(z^{*}).$
The linear projection head $g_{\phi}(h)=\Phi h$ with $\Phi\in\mathbb{R}^{D_{1}\times D_{2}}$ is lifted patch-wise to the block-diagonal operator
$\widetilde{\Phi}:=I_{N}\otimes\Phi\in\mathbb{R}^{ND_{1}\times ND_{2}},$
where $\otimes$ denotes the Kronecker product and $I_{N}$ is the $N\times N$ identity matrix. Since DINOv2 features are $\ell_{2}$–normalized, maximizing the REPA cosine similarity is equivalent to minimizing a squared $\ell_{2}$ distance in the feature space. Thus, the REPA update step applied to $z_{t}$ can be written as
$z_{t}^{(\mathrm{repa})}=z_{t}-\lambda\,\nabla_{z_{t}}\Bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\Bigr\|_{2}^{2}.$
The corresponding diffusion representation is
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{G}_{2}\!\left(z_{t}-\lambda\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}\right).$
Applying a first–order Taylor expansion of $\widetilde{G}_{2}$ around $z_{t}$ for sufficiently small $\lambda$ yields
$\widetilde{h}_{t}^{(\mathrm{repa})}=\widetilde{h}_{t}-\lambda J_{\widetilde{G}_{2}}(z_{t})\,\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}+O(\lambda^{2}),$ (16)
where $J_{\widetilde{G}_{2}}(z_{t})\in\mathbb{R}^{ND_{2}\times d}$ is the Jacobian of $\widetilde{G}_{2}$ at $z_{t}$.
The inner gradient is, by the chain rule,
$\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\Phi\,\widetilde{G}_{2}(z_{t})\bigr\|_{2}^{2}=-2\,J_{\widetilde{G}_{2}}(z_{t})^{\!\top}\,\widetilde{\Phi}^{\top}\,\Big(\widetilde{f}(\bar{x})-\widetilde{\Phi}\,\widetilde{h}_{t}\Big).$
Substituting this into equation 16 gives

### v2

@@ 6 Conclusions
In this work, we introduced representation alignment as a principled framework for solving inverse problems with diffusion- and flow-based generative models. By aligning internal model representations with features from a pretrained DINOv2 encoder, our method provides a strong semantic prior that guides the reconstruction process toward perceptually meaningful solutions. We further offered theoretical insights into this mechanism, showing that REPA regularization corresponds to minimizing a divergence in the DINOv2 feature space and reduces the discrepancy between the current and the clen latent state. These results elucidate why re alignment can enhance perceptual fidelity in inverse problem settings. Overall, our study demonstrates that representation alignment constitutes a flexible and effective tool for improving generative inverse problem solvers, yielding benefits at inference time and integrate seamlessly with a range of diffusion- and flow-based methods.
Impact Statement
This paper presents work aimed at advancing the field of machine learning, with a focus on improving the reliability of inverse problem solvers through representation alignment in pretrained diffusion and flow based models. While this work may have broader societal implications, we do not identify any that require specific discussion here.
Appendix A Appendix.
@@ A.1 Proofs
Proposition A.1.
Assume $x\sim p_{X}$ is a ground-truth image and $\hat{x}\sim p_{\hat{X}\mid Y}$ is its corresponding reconstruction, where $p_{\hat{X}\mid Y}$ is the distribution induced by any reconstruction method. Assume further that all DINOv2 feature embeddings are $\ell_{2}$–normalized. Let $\bar{x}$ be a proxy approximation of $x$ Define the empirical mean DINOv2 feature embedding as
$\mu_{f}(x):=\mathbb{E}_{n}\!\left[f_{\mathrm{DINOv2}}^{[n]}(x)\right].$
Then the expected REPA alignment satisfies
@@
$),\sum_{n=1}^{N}\big\|f_{\text{DINOv2}}^{[n]}(\bar{x})-g_{\phi}(\textsc{DiffEnc}^{[n]}(\hat{x}))\big\|_{2}^{2}=2N\big(1-\text{REPA}(\bar{x},\hat{x}$$\big)$,
@@
Assumption A.2 (Regularity of the diffusion encoder and REPA projection).
We operate under the following regularity and conditioning assumptions on the diffusion feature extractor and the REPA projection. For notational simplicity in the analysis, we introduce an operator corresponding to the vectorized diffusion encoder. Let $\widetilde{G}(\cdot):=\mathrm{vec}(\textsc{DiffEnc}(\cdot,t)):\mathbb{R}^{d}\to\mathbb{R}^{ND}$ denote the vectorized diffusion feature extractor at timestep $t$, and let $\widetilde{g}_{\phi}:\mathbb{R}^{ND}\to\mathbb{R}^{ND}$ denote the patchwise lift of the projection head $g_{\phi}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. Fix two diffusion states $z_{t},z^{*}\in\mathbb{R}^{d}$ and define $u_{t}:=\widetilde{G}(z_{t})$ and $u^{*}:=\widetilde{G}(z^{*})$. Let
$B:=\{z^{*}+s(z_{t}-z^{*}):s\in[0,1]\}\quad\text{and}\quad\mathcal{U}:=\{\widetilde{G}(z):z\in B\}.$
  1. A.2.1
    The mapping $\widetilde{G}$ is differentiable, and its Jacobian $J_{\widetilde{G}}(z)\in\mathbb{R}^{(ND)\times d}$ is Lipschitz continuous on $B$ with Lipschitz constant $L_{G}$.
  2. A.2.2
    There exist constants $m_{G}>0$ and $M_{G}<\infty$ such that
    $J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{G}}(z_{t})\succeq m_{G}^{2}I_{d},\qquad\|J_{\widetilde{G}}(z)\|_{2}\leq M_{G}\quad\text{for all }z\in B.$
  3. A.2.3
    The mapping $\widetilde{g}_{\phi}$ is differentiable, and its Jacobian $J_{\widetilde{g}_{\phi}}(u)\in\mathbb{R}^{(ND)\times(ND)}$ is uniformly well-conditioned at $u_{t}$: there exists a constant $m_{\phi}>0$ such that
    $J_{\widetilde{g}_{\phi}}(u_{t})^{\top}J_{\widetilde{g}_{\phi}}(u_{t})\succeq m_{\phi}^{2}I_{ND},\qquad\|J_{\widetilde{g}_{\phi}}(u)\|_{2}\leq M_{\phi}\quad\text{for all }u\in\mathcal{U}.$
  4. A.2.4
    The Jacobian $J_{\widetilde{g}_{\phi}}$ is Lipschitz continuous on $\mathcal{U}$ with Lipschitz constant $L_{\phi}$.
Proposition A.3 (Local contraction of REPA in the diffusion state space).
Consider an inverse problem in which an observation $y\in\mathbb{R}^{m}$ is generated from a ground-truth image $x\sim p_{X}$, and let $\bar{x}\sim p_{\bar{X}}$ denote the proxy approximation used during REPA. Let $z_{t}\in\mathbb{R}^{d}$ be the diffusion state at timestep $t$ produced by Algorithm 1, and let $z_{t}^{(\mathrm{REPA})}$ denote the diffusion state after applying the REPA regularization update. Let $z^{*}\in\mathbb{R}^{d}$ denote the latent variable associated with the clean image $x$. Under Assumptions $A_{1}$$A_{4}$, define
$r:=\frac{2m^{2}_{G}\,m^{2}_{\phi}}{L_{\phi}M^{3}_{G}M_{\phi}+L_{G}M^{2}_{\phi}M_{G}.}$
If $\|z_{t}-z^{*}\|_{2}<r$, then there exists a stepsize
$0<\lambda\leq\frac{1}{2M_{G}^{2}\,M_{\phi}^{2}}$
such that the REPA update induces a local contraction in the diffusion state space. In particular,
$\|z_{t}^{(\mathrm{REPA})}-z^{*}\|_{2}\;\leq\;C_{1}\,\|z_{t}-z^{*}\|_{2}+C_{2}\left(\sqrt{\mathrm{ApproxErr}(x,\bar{x})}+\sqrt{\mathrm{MisREPA}(x)}\right),$
where
$C_{1}:=1-\lambda m_{G}m_{\phi}\in(0,1),\qquad C_{2}:=2\,\lambda\,M_{G}\,M_{\phi}\,\sqrt{N}.$
Proof.
We first rewrite all quantities in vectorized form. Let $\mathrm{vec}(\cdot)$ denote the operator that stacks the rows of a matrix into a single vector. For any matrix $H\in\mathbb{R}^{k_{1}\times k_{2}}$, define
@@
For the DINOv2 features $f_{\mathrm{DINOv2}}(x)\in\mathbb{R}^{N\times D}$ we write
$\widetilde{f}(x):=\mathrm{vec}\!\bigl(f_{\mathrm{DINOv2}}(x)\bigr)\in\mathbb{R}^{ND}.$
@@
For the diffusion-model feature extractor $\textsc{DiffEnc}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{N\times D}$, we define its vectorized form by
$\widetilde{G}(z_{t}):=\mathrm{vec}\!\bigl(\textsc{DiffEnc}(z_{t},t)\bigr)\in\mathbb{R}^{ND}.$
Recall that $\phi:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is differentiable. Its lifted Jacobian acts patchwise and satisfies
$J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))=I_{N}\otimes J_{\phi}(G_{2}(z_{t})).$
The REPA update is given by
$z_{t}^{(\mathrm{REPA})}=z_{t}-\lambda\nabla_{z_{t}}\bigl\|\widetilde{f}(\bar{x})-\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))\bigr\|_{2}^{2}.$
By the chain rule,
$z_{t}^{(\mathrm{REPA})}=z_{t}+2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))\bigr).$
Subtracting $z^{*}$ from both sides yields
$z_{t}^{(\mathrm{REPA})}-z^{*}=z_{t}-z^{*}+2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))\bigr).$
We now add and subtract $\widetilde{f}(x)$ and $\widetilde{g}_{\phi}(\widetilde{G}(z^{*}))$ inside the parentheses:
$z_{t}^{(\mathrm{REPA})}-z^{*}$ $=(z_{t}-z^{*})-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))-\widetilde{g}_{\phi}(\widetilde{G}(z^{*}))\bigr)$
$\quad+2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{f}(\bar{x})-\widetilde{f}(x)\bigr)$
$\quad+2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{f}(x)-\widetilde{g}_{\phi}(\widetilde{G}(z^{*}))\bigr).$ (12)
We define the contractive part as
$\mathcal{T}_{t}:=(z_{t}-z^{*})-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\bigl(\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))-\widetilde{g}_{\phi}(\widetilde{G}(z^{*}))\bigr).$
Applying the fundamental theorem of calculus to $u\mapsto\widetilde{g}_{\phi}(u)$ along $u(s)=\widetilde{G}(z(s))$ with $z(s)=z^{*}+s(z_{t}-z^{*})$, $s\in[0,1]$, yields
$\widetilde{g}_{\phi}(\widetilde{G}(z_{t}))-\widetilde{g}_{\phi}(\widetilde{G}(z^{*}))=\int_{0}^{1}J_{\widetilde{g}_{\phi}}\!\bigl(\widetilde{G}(z(s))\bigr)\,J_{\widetilde{G}}\!\bigl(z(s)\bigr)\,(z_{t}-z^{*})\,ds.$
Substituting into the definition of $\mathcal{T}_{t}$ gives
$\mathcal{T}_{t}$ $=\Bigl[I-2\lambda\int_{0}^{1}J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}\!\bigl(\widetilde{G}(z_{t})\bigr)^{\top}J_{\widetilde{g}_{\phi}}\!\bigl(\widetilde{G}(z(s))\bigr)J_{\widetilde{G}}\!\bigl(z(s)\bigr)\,ds\Bigr](z_{t}-z^{*}).$ (13)
Add and subtract $J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})$ inside the integral to obtain
$\mathcal{T}_{t}$ $=\Bigl[I-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))J_{\widetilde{G}}(z_{t})\Bigr](z_{t}-z^{*})$
$\quad-2\lambda\Bigl[\int_{0}^{1}J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\Bigl(J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr)\,ds\Bigr](z_{t}-z^{*}).$ (14)
Taking Euclidean norms in (14) and using the triangle inequality gives
$\|\mathcal{T}_{t}\|_{2}$ $\leq\Bigl\|I-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\;\|z_{t}-z^{*}\|_{2}$
$\quad+2\lambda\Bigl\|\int_{0}^{1}J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\Bigl(J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr)\,ds\Bigr\|_{2}\;\|z_{t}-z^{*}\|_{2}.$ (15)
Moreover, by submultiplicativity of the spectral norm and $\bigl\|\int_{0}^{1}A(s)\,ds\bigr\|_{2}\leq\int_{0}^{1}\|A(s)\|_{2}\,ds$, the second term can be bounded as
$\Bigl\|$ $\int_{0}^{1}J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}\Bigl(J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr)\,ds\Bigr\|_{2}\leq$ (16)
$\int_{0}^{1}\|J_{\widetilde{G}}(z_{t})\|_{2}\;\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\Bigl\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\,ds.$ (17)
Combining (15)–(17) yields
$\|\mathcal{T}_{t}\|_{2}\leq\Bigl\|I-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\;\|z_{t}-z^{*}\|_{2}$
$\quad+2\lambda\Bigl(\int_{0}^{1}\|J_{\widetilde{G}}(z_{t})\|_{2}\;\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\Bigl\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\,ds\Bigr)\|z_{t}-z^{*}\|_{2}.$ (18)
We bound the difference inside the integral by writing
$\Bigl\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr\|_{2}$
$\leq\Bigl\|\bigl(J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\bigr)J_{\widetilde{G}}(z(s))\Bigr\|_{2}+\Bigl\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\bigl(J_{\widetilde{G}}(z(s))-J_{\widetilde{G}}(z_{t})\bigr)\Bigr\|_{2}$
$\leq\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\;\|J_{\widetilde{G}}(z(s))\|_{2}+\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\;\|J_{\widetilde{G}}(z(s))-J_{\widetilde{G}}(z_{t})\|_{2}.$ (19)
Using the Lipschitz property of $J_{\widetilde{G}}$,
$\|J_{\widetilde{G}}(z(s))-J_{\widetilde{G}}(z_{t})\|_{2}\leq L_{G}\,\|z(s)-z_{t}\|_{2}=L_{G}(1-s)\,\|z_{t}-z^{*}\|_{2},$
together with the Lipschitz continuity of $J_{\widetilde{g}_{\phi}}$,
$\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\leq L_{\phi}M_{G}(1-s)\,\|z_{t}-z^{*}\|_{2},$
and the uniform bounds
$\|J_{\widetilde{G}}(z(s))\|_{2}\leq M_{G},\qquad\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\leq M_{\phi},$
This yields
$\Bigl\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z(s)))\,J_{\widetilde{G}}(z(s))-J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\,J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\leq(1-s)\Bigl(L_{\phi}M^{2}_{G}+L_{G}M_{\phi}\Bigr)\|z_{t}-z^{*}\|_{2}.$
Substituting this bound into (18) and integrating over $s\in[0,1]$ yields
$\|\mathcal{T}_{t}\|_{2}$ $\leq\Bigl\|I-2\lambda\,J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))J_{\widetilde{G}}(z_{t})\Bigr\|_{2}\;\|z_{t}-z^{*}\|_{2}$
$\quad+\lambda\Bigl(L_{\phi}M^{3}_{G}M_{\phi}+L_{G}M^{2}_{\phi}M_{G}\Bigr)\|z_{t}-z^{*}\|_{2}^{2}.$ (20)
Since the symmetric matrix
$S_{t}:=J_{\widetilde{G}}(z_{t})^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))^{\top}J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))J_{\widetilde{G}}(z_{t})$
has eigenvalues contained in $m^{2}_{G}m^{2}_{\phi},\,M^{2}_{G}M^{2}_{\phi}]$, its spectral norm satisfies
$\bigl\|I-2\lambda S_{t}\bigr\|_{2}=\max_{\sigma\in[m^{2}_{G}m^{2}_{\phi},\,M^{2}_{G}M^{2}_{\phi}]}|1-2\lambda\sigma|.$
For any $0<\lambda\leq 1/(2M^{2}_{G}M^{2}_{\phi})$, the maximum is attained at $\sigma=m^{2}_{G}m^{2}_{\phi}$, and hence
$\bigl\|I-2\lambda S_{t}\bigr\|_{2}\leq 1-2\lambda m^{2}_{G}m^{2}_{\phi}.$
Substituting this bound into (20) yields
$\|\mathcal{T}_{t}\|_{2}$ $\leq\Bigl(1-2\lambda m^{2}_{G}m^{2}_{\phi}+\lambda\Bigl(L_{\phi}L_{G}M^{2}_{G}M_{\phi}+L_{G}M^{2}_{\phi}M_{G}\Bigr)\Bigr)\|z_{t}-z^{*}\|_{2}\Bigr)\|z_{t}-z^{*}\|_{2}.$ (21)
Therefore, if
$\|z_{t}-z^{*}\|_{2}<\frac{2m^{2}_{G}\,m^{2}_{\phi}}{L_{\phi}M^{3}_{G}M_{\phi}+L_{G}M^{2}_{\phi}M_{G}\ }\qquad\text{and}\qquad 0<\lambda\leq\frac{1}{2M_{G}^{2}M_{\phi}^{2}},$
then the contractive part satisfies the strict contraction
$\|\mathcal{T}_{t}\|_{2}\leq\bigl(1-2\lambda m_{G}^{2}m_{\phi}^{2}\bigr)\,\|z_{t}-z^{*}\|_{2},$
with contraction factor strictly smaller than one. By the definitions of the approximation and misalignment errors,
$\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}=\sqrt{N}\,\sqrt{\mathrm{ApproxErr}(x,\bar{x})},\qquad\|\widetilde{f}(x)-\widetilde{g}_{\phi}\,(\widetilde{G}(z^{*}))\|_{2}=\sqrt{N}\,\sqrt{\mathrm{MisREPA}(x)}.$
Returning to the decomposition we combine the contraction bound for $\mathcal{T}_{t}$ with the remaining two terms. Using the triangle inequality together with the uniform bounds $\|J_{\widetilde{G}}(z_{t})\|_{2}\leq M_{G}$ and $\|J_{\widetilde{g}_{\phi}}(\widetilde{G}(z_{t}))\|_{2}\leq M_{\phi}$, we obtain
$\|z_{t}^{(\mathrm{repa})}-z^{*}\|_{2}$ $\leq\|\mathcal{T}_{t}\|_{2}+2\lambda M_{G}M_{\phi}\,\|\widetilde{f}(\bar{x})-\widetilde{f}(x)\|_{2}$
$\quad+2\lambda M_{G}M_{\phi}\,\|\widetilde{f}(x)-\widetilde{g}_{\phi}(\,\widetilde{G}(z^{*}))\|_{2}.$ (22)
Substituting the contraction estimate together with the above identities yields